OpenStudy (goformit100):

If y=| cos x - sin x |, find f'(90 degree)

OpenStudy (goformit100):

@zepdrix

zepdrix (zepdrix):

This one is a little annoying. You can either remember this little fact:\[\Large\bf\sf \frac{d}{dx}|x|\quad=\quad \frac{x}{|x|}\]Or we can work it out the long way if necessary.

OpenStudy (goformit100):

please provide me the long way...... I want to learn it from....the longer one

zepdrix (zepdrix):

\[\Large\bf\sf |x|\quad:=\sqrt{x^2}\]So we'll start with this definition and take the derivative,

zepdrix (zepdrix):

\[\Large\bf\sf \left(\sqrt {x^2}\right)'\quad=\quad\frac{1}{2\sqrt{x^2}}\cdot(x^2)'\]

zepdrix (zepdrix):

Understand what I did there? :o Do you remember the derivative of sqrt(x) ? It's a good one to have memorized.

OpenStudy (goformit100):

yes got it till here, GO ON

zepdrix (zepdrix):

So if we cancel the 2's out,\[\Large\bf\sf \left(\sqrt {x^2}\right)'\quad=\quad\frac{x}{\sqrt{x^2}}\]And recall our denominator was defined to be |x| before, so we'll put it back in now.\[\Large\bf\sf |x|'\quad=\quad\frac{x}{|x|}\]

zepdrix (zepdrix):

Mmm I guess that wasn't super long XD lol. With this specific problem you have though, it might be a little weird when we go to evaluate it at 90 degrees. We'll need to be careful.

OpenStudy (goformit100):

ok then do it in radian .... I can convert it

OpenStudy (anonymous):

1) find f(90 degrees) f(90 degrees) = |cos(x)-sin(x)|=sin(x)-cos(x)=1 2) take the derivative of sin(x)-cos(x) 3) plug in for x=90 degrees

OpenStudy (goformit100):

Explain this step \Large\bf\sf |x|\quad:=\sqrt{x^2}

OpenStudy (goformit100):

How to use the theory taught by you here ?

zepdrix (zepdrix):

Ummm I like eashy's method a little better. You have less stuff floating around. At 90 degrees, our function |cos x-sin x| = -(cos x-sin x) = sin x - cos x. And it's much easier to take a derivative from there. The way I was doing it, we would apply the derivative,\[\Large\bf\sf |\cos x-\sin x|'\quad=\quad \frac{\cos x - \sin x}{|\cos x - \sin x|}\cdot \left(\cos x - \sin x\right)'\]We get the derivative of the absolute, but then we have to apply the chain rule. This should still work, it's just a bit burdensome with all of these pieces.

OpenStudy (goformit100):

Please explain this step |cos x-sin x| = -(cos x-sin x) = sin x - cos x

zepdrix (zepdrix):

Here is another way we can define absolute value, as a piece-wise function,\[\Large |x|=\cases{\quad x,\quad x>0\\-x, \quad x<0}\] So applying this to our problem,\[\Large |\cos x-\sin x|=\cases{\quad \cos x-\sin x,\qquad \cos x-\sin x>0\\-(\cos x-\sin x), \quad \cos x- \sin x<0}\]

zepdrix (zepdrix):

At x=90 degrees,\[\Large\bf\sf y(90^o)\quad=\quad |\cos90^o-\sin90^o|\quad=\quad |0-1|\]See how the inside is negative? That means our absolute function is negative at this particular x coordinate,\[\Large |\cos x-\sin x|=\cases{\quad \cos x-\sin x,\qquad \cos x-\sin x>0\\\color{orangered}{-(\cos x-\sin x), \quad \cos x- \sin x<0}}\]The cosx - sinx gave us something less than 0. So we use this part of our piece-wise function.

zepdrix (zepdrix):

Sorry if this is super confusing :( Absolute was always a bit weird to me also.. Trying to figure out the right way to word this.

OpenStudy (goformit100):

how you thought that cos x - sin x is < 0

zepdrix (zepdrix):

\[\Large\bf\sf y(x)=|\color{royalblue}{\cos x-\sin x}|\]The blue part is what tells us what sign to use on our absolute when we want to drop the bars. So for example:\[\Large\bf\sf y=|x|\]At x=-2,\[\Large\bf\sf y=|-2|\]You might think, "Oh the absolute bars tell us to destroy the negative sign and just give us a positive." But that's not what's happening. If we remove the absolute bars we can maybe more clearly see what's going on. \[\Large\bf\sf y=(-2)\]The absolute is telling us that this needs to be positive. So we need a negative in front to make it positive,\[\Large\bf\sf y=-(-2)\] So the example is telling us that at x=-2, our function \(\Large\bf\sf y=|x|\) is actually \(\Large\bf\sf y=-x\)

zepdrix (zepdrix):

So to answer your question, ummm...

zepdrix (zepdrix):

\[\Large\bf\sf y(x)=|\color{royalblue}{\cos x-\sin x}|\]When the blue stuff is negative, our function is actually:\[\Large\bf\sf y(x)=-(\cos x-\sin x)\]

zepdrix (zepdrix):

That's what's getting rid of the negative value inside of the bars.

OpenStudy (goformit100):

okay

OpenStudy (goformit100):

but why particularly this | cos x - sin x |

zepdrix (zepdrix):

Hmm? 0_o

zepdrix (zepdrix):

At x=90 degrees, (the stuff in the brackets is negative. So we put a negative in front to compensate. Our function is,\[\Large\bf\sf y(x)\quad=\quad \sin x- \cos x\]I distributed the negative sign.

zepdrix (zepdrix):

And it's easier to take a derivative from there.

OpenStudy (goformit100):

Thanks sir